# Deconfined quantum critical points in fermionic systems with spin-charge separation

### v3

@@ II.2 Beyond Landau phase transitions
In order to better exploit this crucial point, it is convenient to first relate the decay of correlation functions at the transition points to the Luttinger parameters $K_{\nu}$. In this regards, see A for more details, bosonization shows $K_{\nu}\approx 0$ in presence of $\Delta_{\nu}\neq 0$ while $K_{\nu}$ remains thermodynamically finite in presence of a vanishing gap in the $\nu$ channel. This point results quite important as the Tomonaga Luttinger theory Tomonaga (1950); Luttinger (1963) predicts the $K_{\nu}$’s to induce in the MI, LEL, and LL phases different algebraic decay of the correlation functions $C_{X}(r)$ (with $X=CDW,AF,BOW$) having long-range order in the $X$ LO phase. Notably, these latter can have the same decay only in a fully gapless Luttinger liquid
@@
holds. In addition, the fixed value $\varphi_{s}=0$ suggests that long-range of the spin parity correlation $C_{P}^{s}(r)=\langle O^{s\dagger}_{P}(x)O_{P}^{s}(x+r)\rangle$ has to occur. Analogously, at the BOW-CDW transition $K_{c}=0$, therefore
@@ III Microscopic Hamiltonian
In addition, the competition between intersite and onsite interaction of the form of
@@
can give rise to BOW ordering Nakamura (2000). Finally, away from strongly interacting perturbative regimes, quantum fluctuations are induced by tunneling process
@@
Notably, this latter presents strong connections with the Hamiltonians Julià-Farré et al. (2022); Guardado-Sanchez et al. (2021); Su et al. (2023); Carroll et al. (2024) under investigations in the context of ultracold atomic systems. As these setups are characterized by an unprecedented level of accuracy and versatility, here we restrict our analysis to the regime of half-filling and vanishing total magnetization, i.e. $\sum_{j}S^{\nu}_{j}=0$ for both $\nu=c,s$.
@@ III.1 Bosonization analysis
Thus along such line, a continuous CDW-AF transition is predicted, in principle compatible with a fully gapless DQCP. However, it is a well known fact that when the above analysis is complemented by other results, for instance by level crossing analysis Nakamura (2000), the actual transition lines observed in the two channels split: the BOW-CDW and BOW-AF transitions will occur in the weak coupling region, while they will merge into the CDW-AF transition at intermediate couplings only. If this is the case, in the weakly coupled regime this theory turns out to be compatible with the presence of the partly gapped DQCPs envisaged previously.
@@ III.2 VUMPS analysis
Since bosonization is believed to provide accurate results in the low energy limit, in the following we employ numerical calculations to explore the possible appearance of DQCPs away from this regime. To unveil this point, we perform a numerical analysis based on the VUMPS technique. Here, the state is described by an infinitely repeating unit cell of tensors defined on the sites of the lattice. The tensors are then variationally optimized with a procedure similar to DMRG White (1992), by using the rest of the (infinite) system as an environment. Compared to infinite DMRG, VUMPS has the advantage to completely replace the full environment at each step, making its convergence significantly faster. This is crucial especially at critical points, where a larger bond dimension is required to get accurate results. Moreover, the absence of boundaries in the variational state makes it easier to escape metastable states that can become more favourable around the critical point where two phases with different symmetries meet. This being exactly the situation that we find at the phase transitions we are studying. Indeed, the use of VUMPS has proved critical in the analysis of DQCPs in other systems Roberts et al. (2019); Baldelli et al. (2024).
@@ III.2.1 BOW-CDW phase transition
As shown in Fig. (1)(a), for $J_{z}=0$, the Hamiltonian in Eq.(16) is the known extended Fermi-Hubbard model Nakamura (2000); Sengupta et al. (2002); Sandvik et al. (2004); Ejima and Nishimoto (2007); Zhang (2004). For dominating $U/t$, the system is in MI.
@@
The situation changes drastically for larger values of $V$. Here, two LO phases can take place: a BOW phase when $V\sim U$ are away from strongly interacting perturbative regimes and a CDW for dominant intersite repulsion.
@@
As discussed in the previous sections, this quantity is expected to be zero for a finite $\Delta_{\nu}$, and finite in the case of a gapless $\nu$-channel. To perform the finite-size scaling, we construct finite MPS of length $L$ by repeating $L/2$ unit cells obtained from the VUMPS optimization. Although our analysis considers relatively large interactions $U/t=4$, our VUMPS simulations provide results compatible with the field theory. On one hand, in Fig.2(a) we indeed find that $K_{s}\approx 0$ at the CDW-BOW transition point, meaning that $\Delta_{s}\neq 0$.
@@
On the other, $K_{c}$ shows a $L$-dependence compatible with a vanishing thermodynamic value for all the considered values of $V$, except at the transition point where the size dependence is absent. This strongly points in the direction that a continuous phase transition between the two LO regimes persists not only in low energy limit. In order to enforce this aspect, we extract the correlation length $\xi_{\nu}(\chi)$ relative to the $\nu$ degrees of freedom  [^2] The correlation length is extracted through the relation $\xi_{\nu}=-N/\log(\lambda_{2})$ where $N$ is the number of sites of a unit cell and $\lambda_{2}$ is second highest eigenvalue of the transfer matrix.. A genuine second-order phase transition can indeed be identified by considering the scaling of the correlation length $\xi_{\nu}$ as a function of the bond dimension $\chi$. In particular, one has to expect that in presence of a single point gap closing, the correlation length would show a $\chi$-dependence in the form of a cusp uniquely at that point. Fig. 2(b) shows this behavior for $\xi_{c}$, thus unambiguously confirming the continuous nature of this transition. Notably, this charge gap closing clearly supports the phenomenon of symmetry enriching. Indeed, based on the previous symmetry arguments, the presence of a gapless sector is associated to a $SU(2)$ symmetry. Therefore, it is straightforward to realize that an emergent $SU(2)$ symmetry relative to the charge degrees of freedom which is not present either in BOW and CDW takes place at the CDW-BOW transition point. Further relevant informations can be derived by calculating the decay of the correlators in $C_{CDW}(r)$ and $C_{BOW}(r)$. Fig. 2(c) clearly proves their expected algebraic decay at the transition point. In addition, the same figure confirms that the spin parity parity correlator has long-range order, i.e. $C_{P}^{s}(r)\neq 0$ for $r\rightarrow\infty$.
Based on this analysis, we have strong indications that peculiar DQCPs can occur in one-dimensional interacting fermionic systems. As underlined, previous examples found indeed fully gapless DQCPs. As a consequence of the phenomenon of spin-charge separation, our analysis shows instead strong evidences on the presence DQCPs with a finite gap and long-range order of a nonlocal order parameter.
@@
where $V_{C}$ is the critical value of the intersite interaction where we detect the charge gap closing. Crucially, while an Ising phase transition would formally imply $\beta_{BOW/CDW}=1/8$, our results based on an extrapolation in the limit $\chi\rightarrow\infty$, see B, find different values of such critical exponent thus proving this phase transition as a novel example of a 1D DQCP. As pointed out, with respect to previously investigated DQCPs here such beyond Landau’s transition points are further associated to a gapped excitation spectrum giving rise to long-range order of the spin parity operator.
Notably, in agreement with previous finite-size studies on the same model  Ejima and Nishimoto (2007), we find the BOW-CDW phase transition to become discontinuous for larger values of the interactions and therefore the partially gapped DQCPs disappear. However, this result is certainly not surprising. When increasing the contact interaction $U$, the mean-field treatment of the transition associated to negligible quantum fluctuations becomes indeed more accurate and a genuine first-order transition is expected. We prove this point for $U=8t$. In this case, the order parameters $\mathcal{O}_{BOW}$ and $\mathcal{O}_{CDW}$ exhibit a sharp jump at the critical point independent of the bond dimension $\chi$, see Fig. 4(a), thus discarding the possibility of a single point gap closing. In addition, the Luttinger parameters $K_{\nu}$ reported in Fig. 4(b) show both a clear discontinuity a the transition point and very small size dependent values thus compatible with a thermodynamic vanishing behaviour.
@@ III.2.2 AF-BOW phase transition
We now move to the analysis of the model in Eq. (16) for $J_{z}>0$ which phase diagram is shown in Fig. (1)(b). As specified, a finite antiferromagnetic coupling allows for the breaking of the discrete spin flip symmetry. A part from slightly moving the transition points, we find indeed that the main role played by small $J_{z}$ is to replace in the Hamiltonian ground state the disordered MI with a locally ordered AF phase. Thanks to this aspect, we can now perform an analysis analogous to the one made in the previous section to investigate whether the AF-BOW transition point can also be classified as a partially gapped DQCP. Here, we focus on intermediate interactions $U/t=4$ and $J_{z}/t=1$ while varying $V/t$ and we again start by exploring the behavior of the Luttinger parameters. Our results in Fig. 5(a) find that at the transition point $K_{c}\approx 0$ 333the increasing of $K_{c}$ for larger values of $V/t$ is due to the fact that the system is approaching the BOW-CDW transition while $K_{s}$ shows a finite value not affected by different $L$. This analysis combined with the clear cusp of $\xi_{s}(\chi)$ shown in Fig. 5(b) demonstrates that the AF-BOW is also a continuous phase transition where the spin gap closes and $\Delta_{c}$ remains finite. Based of the already pointed out symmetry arguments, this results proves the emergence at the critical point of a $SU(2)$ symmetry related to the spin degrees of freedom not present in either the BOW or the AF phase. Fig. 5(c) also demonstrates that such symmetry enriched transition point is associated to long-range order of the $C_{P}^{c}$ while the correlators $C_{AF}$ and $C_{BOW}$ decay algebraically. Thanks to these findings, we can again infer that the phase transition between the two locally ordered AF and BOW phases is continuous and thus supporting the possible appearance of a partly gapped DQCP. The results in Fig. 6 confirm this intuition. As before, we can indeed show that local order parameters $\mathcal{O}_{AF}$ and $\mathcal{O}_{BOW}$ vanish continuously at the transition point. Moreover, through a $\chi$-extrapolation we calculate their critical exponent
@@
and we find $\beta_{AF}\sim 0.525$ and $\beta_{BOW}\sim 0.46$, again discarding the possibility of an Ising transition. As a consequence, this result proves again that the decoupling between spin an charge degrees of freedom can support the appearance a partly gapped DQCP which in the specific BOW-AF case is characterized by $\Delta_{c}\neq 0$ while having long-range order of $C_{P}^{c}(r)$.
@@ III.2.3 AF-CDW phase transition
Finally, we study the AF-CDW phase transition. As already pointed out at the level of field theory, in order to find such transition to be continuous, both gaps have to vanish at the transition point. This aspect clearly rules out the possible presence of a partly gapped DQCPs while, in principle, it still allows for standard fully gapless DQCPs where all the correlators decay algebraically. As already discussed, the presence of such a DQCPs would require a fine tuning of the microscopic parameters in order to have a simultaneous closing of both the spin and charge gaps. In this regard, our numerical analysis finds that the AF-CDW transition is always first order for all values of $U/t$. In Fig. 7 we prove this for the case of $U/t=8$. As can be seen, we find indeed that both the order parameters ${\cal O}_{CDW}$ ${\cal O}_{AF}$ and the Luttinger parameters display a discontinuity at the transition point, see Fig. 7(a) and (b) respectively. Moreover, while the former quantities are not affected by different $\chi$ values, the latter always show size dependence associated to almost vanishing values.
@@ IV Conclusions
We unveiled a new type of deconfined quantum critical points. Contrary to previously known examples where they are associated to fully gapless excitations, we have shown that the phenomenon of spin and charge separation can generate partly gapped deconfined quantum critical points. Specifically, by means of analytical and numerical treatments, we have been able to prove that specific phase transitions connecting locally ordered fermionic phases are not captured by the Landau-Ginzburg-Wilson symmetry-breaking paradigm. Notably, at these phase transitions we find that only one gap closes while the other remains finite. Therefore, long-range order of specific nonlocal correlation functions occurs. In addition, we calculated the critical exponents governing how the local order parameters vanish. In such a way, we demonstrated that these transitions are one-dimensional realizations of deconfined quantum critical points with the intriguing feature of being only partly gapped. In conclusion, we also point out that at least one of these phase transitions, the one between the bond-order-wave and charge-density-wave, has been predicted to characterize the phase diagram of ultracold dipolar atoms in optical lattice Di Dio et al. (2014); Julià-Farré et al. (2022). Here, the presence of partially gapped deconfined quantum critical points can be probed by measuring the defect production Keesling et al. (2019); Xi and Yu (2022); Shu et al. (2022) predicted by the Kibble-Zurek mechanism Kibble (1976, 1980); Zurek (1985, 1996); Polkovnikov et al. (2011); del Campo and Zurek (2014) in continuous transitions. In addition, quantum gas microscopy Gross and Bakr (2021) and noise correlators measurements Gallego-Lizarribar et al. (2024) allow for an accurate detection of local density Endres et al. (2011) and spin Parsons et al. (2016) orderings and therefore for an accurate probing of nonlocal order parameters and locally ordered phases. As consequence, we believe that our results not only unveil a novel interesting type of deconfined quantum critical points but also pave the way towards their experimental detection, which represents a fundamental challenge solid state platforms.  Zayed et al. (2017); Guo et al. (2020); Hong et al. (2022); Cui et al. (2023).

### v4

@@ II.2 Beyond Landau phase transitions
In order to better exploit this crucial point, it is convenient to first relate the decay of correlation functions at the transition points to the Luttinger parameters $K_{\nu}$. In this regards, see A for more details, bosonization shows $K_{\nu}\approx 0$ in presence of $\Delta_{\nu}\neq 0$ while $K_{\nu}$ remains thermodynamically finite in presence of a vanishing gap in the $\nu$ channel. This point is quite important as the Tomonaga Luttinger theory [45, 46] predicts the $K_{\nu}$’s to induce in the MI, LEL, and LL phases different algebraic decay of the correlation functions $C_{X}(r)$ (with $X=CDW,AF,BOW$) having long-range order in the $X$ LO phase. Notably, these latter can have the same decay only in a fully gapless Luttinger liquid
@@
holds. In addition, the fixed value $\varphi_{s}=0$ suggests that long-range order of the spin parity correlation $C_{P}^{s}(r)=\langle O^{s\dagger}_{P}(x)O_{P}^{s}(x+r)\rangle$ has to occur. Analogously, at the BOW-AF transition $K_{c}=0$, therefore
@@ III Microscopic Hamiltonian
Here, the operators $S_{j}^{c}$ and $S^{s}_{j}$ are the ones defined in Eq.(3). In addition, the competition between intersite and onsite interaction of the form of
@@
can give rise to BOW ordering [82]. Finally, away from strongly interacting perturbative regimes, quantum fluctuations are induced by tunneling process
@@
Notably, this latter presents strong connections with the Hamiltonians [78, 83, 84, 85] under investigations in the context of ultracold atomic systems. As these setups are characterized by an unprecedented level of accuracy and versatility, here we restrict our analysis to the regime of half-filling and vanishing total magnetization, i.e. $\sum_{j}S^{\nu}_{j}=0$ for both $\nu=c,s$.
@@ III.1 Bosonization analysis
Thus, along such line, a continuous CDW-AF transition is predicted, in principle compatible with a fully gapless DQCP. However, it is a well known fact that when the above analysis is complemented by other results, for instance by level crossing analysis [82], the actual transition lines observed in the two channels split: the BOW-CDW and BOW-AF transitions will occur in the weak coupling region, while they will merge into the CDW-AF transition at intermediate couplings only. If this is the case, in the weakly coupled regime this theory turns out to be compatible with the presence of the partially gapped DQCPs envisaged previously.
@@ III.2 VUMPS analysis
Since bosonization is believed to provide accurate results in the low energy limit, in the following we employ numerical calculations to explore the possible appearance of DQCPs away from this regime. To unveil this point, we perform a numerical analysis based on the VUMPS technique. Here, the state is described by an infinitely repeating unit cell of tensors defined on the sites of the lattice. The tensors are then variationally optimized with a procedure similar to DMRG [86], by using the rest of the (infinite) system as an environment. Compared to infinite DMRG, VUMPS has the advantage to completely replace the full environment at each step, making its convergence significantly faster. This is crucial especially at critical points, where a larger bond dimension is required to get accurate results. Moreover, the absence of boundaries in the variational state makes it easier to escape metastable states that can become more favourable around the critical points in study, where two states that break different symmetries have comparable energies. Indeed, the use of VUMPS has proved critical in the analysis of DQCPs in other systems [34, 39].
@@ III.2.1 BOW-CDW phase transition
As shown in Fig. (1)(a), for $J_{z}=0$, the Hamiltonian in Eq.(17) is the known extended Fermi-Hubbard model [82, 87, 88, 89, 90]. For dominating $U/t$, the system is in a MI phase.
@@
The situation changes drastically for larger values of $V$. Here, two LO phases can take place: a BOW phase when $V\sim U$ are away from strongly interacting perturbative regimes, and a CDW for dominant intersite repulsion.
@@
As discussed in the previous sections, this quantity is expected to be zero for a finite $\Delta_{\nu}$, and finite in the case of a gapless $\nu$-channel. To perform the finite-size scaling, we construct multiple finite MPS of length $L$ by repeating $L/2$ unit cells obtained from the VUMPS optimization. Although our analysis considers relatively large interactions $U/t=4$, our VUMPS simulations provide results compatible with the field theory. On one hand, in Fig.2(a) we indeed find that $K_{s}\approx 0$ at the CDW-BOW transition point, meaning that $\Delta_{s}\neq 0$.
@@
On the other, $K_{c}$ shows a $L$-dependence compatible with a vanishing thermodynamic value for all the considered values of $V$, except at the transition point where the size dependence is absent. This strongly points in the direction that a continuous phase transition between the two LO regimes persists not only in the low energy limit. In order to enforce this aspect, we extract the correlation length $\xi_{\nu}(\chi)$ relative to the $\nu$ degrees of freedom  [^3] The correlation length is extracted through the relation $\xi_{\nu}=-N/\log(\lambda_{2})$ where $N$ is the number of sites of a unit cell and $\lambda_{2}$ is second highest eigenvalue of the transfer matrix.. A genuine second-order phase transition can indeed be identified by considering the scaling of the correlation length $\xi_{\nu}$ as a function of the bond dimension $\chi$. In particular, one has to expect that in presence of a single point gap closing, the correlation length would show a $\chi$-dependence in the form of a cusp uniquely at that point. Fig. 2(b) shows this behavior for $\xi_{c}$, thus strongly supporting the continuous nature of this transition. Notably, this charge gap closing clearly supports the phenomenon of symmetry enriching. Indeed, based on known symmetry properties of SG model (see Appendix), the presence of a gapless sector is associated to a $U(1)\times U(1)$ (possibly $SU(2)$) symmetry in the same sector. It is straightforward to realize that such emergent symmetry relative to the charge degrees of freedom –not present either in BOW and CDW phases– takes place at the CDW-BOW transition point. Further relevant information can be derived by calculating the decay of the correlators in $C_{CDW}(r)$ and $C_{BOW}(r)$. Fig. 2(c) clearly shows their expected algebraic decay at the transition point. In addition, the same figure confirms that the spin parity parity correlator has long-range order, i.e. $C_{P}^{s}(r)\neq 0$ for $r\rightarrow\infty$.
Based on this analysis, we have strong indications that peculiar DQCPs can occur in one-dimensional interacting fermionic systems. As underlined, previous examples found indeed fully gapless DQCPs. As a consequence of the phenomenon of spin-charge separation, our analysis shows instead strong evidences on the presence of DQCPs with a finite gap and long-range order of a nonlocal order parameter.
@@
where $V_{C}$ is the critical value of the intersite interaction where we detect the charge gap closing. Crucially, while an Ising phase transition would formally imply $\beta_{BOW/CDW}=1/8$, our results based on an extrapolation in the limit $\chi\rightarrow\infty$, see B, find different values of such critical exponent thus proving this phase transition as an example of a 1D DQCP. As pointed out, with respect to previously investigated DQCPs here such beyond Landau’s transition points are further associated to a gapped excitation spectrum giving rise to long-range order of the spin parity operator.
Notably, in agreement with previous finite-size studies on the same model  [89], we find the BOW-CDW phase transition becomes discontinuous for larger values of the interactions and therefore the partially gapped DQCPs disappear. However, this result is certainly not surprising. When increasing the contact interaction $U$, the mean-field treatment of the transition associated to negligible quantum fluctuations becomes indeed more accurate and a genuine first-order transition is expected. We prove this point for $U=8t$. In this case, the order parameters $\mathcal{O}_{BOW}$ and $\mathcal{O}_{CDW}$ exhibit a sharp jump at the critical point independent of the bond dimension $\chi$, see Fig. 4(a), thus discarding the possibility of a single point gap closing. In addition, the Luttinger parameters $K_{\nu}$ reported in Fig. 4(b) show both a clear discontinuity a the transition point and very small size dependent values thus compatible with a thermodynamic vanishing behaviour.
@@ III.2.2 AF-BOW phase transition
We now move to the analysis of the model in Eq. (17) for $J_{z}>0$ which phase diagram is shown in Fig. (1)(b). As specified, a finite antiferromagnetic coupling allows for the breaking of the discrete spin flip symmetry. A part from slightly moving the transition points, we find indeed that the main role played by small $J_{z}$ is to replace in the Hamiltonian ground state the disordered MI with a locally ordered AF phase. Thanks to this aspect, we can now perform an analysis analogous to the one made in the previous section to investigate whether the AF-BOW transition point can also be classified as a partially gapped DQCP. Here, we focus on intermediate interactions $U/t=4$ and $J_{z}/t=1$ while varying $V/t$ and we again start by exploring the behavior of the Luttinger parameters. Our results in Fig. 5(a) find that at the transition point $K_{c}\approx 0$ 444the increasing of $K_{c}$ for larger values of $V/t$ is due to the fact that the system is approaching the BOW-CDW transition while $K_{s}$ shows a finite value not affected by different $L$. This analysis combined with the clear cusp of $\xi_{s}(\chi)$ shown in Fig. 5(b) demonstrates that the AF-BOW is also a continuous phase transition where the spin gap closes and $\Delta_{c}$ remains finite. Based of the already pointed out symmetry arguments, this results proves the emergence at the critical point of a $U(1)\times U(1)$ symmetry in the spin channel not present in either the BOW or the AF phase. Fig. 5(c) also demonstrates that such symmetry enriched transition point is associated to long-range order of the $C_{P}^{c}$ while the correlators $C_{AF}$ and $C_{BOW}$ decay algebraically. Thanks to these findings, we can again infer that the phase transition between the two locally ordered AF and BOW phases is continuous and thus supporting the possible appearance of a partly gapped DQCP. The results in Fig. 6 confirm this intuition. As before, we can indeed show that local order parameters $\mathcal{O}_{AF}$ and $\mathcal{O}_{BOW}$ vanish continuously at the transition point. Moreover, through a $\chi$-extrapolation we calculate their critical exponent
@@
and we find $\beta_{AF}\sim 0.525$ and $\beta_{BOW}\sim 0.46$, again discarding the possibility of an Ising transition. As a consequence, this result proves again that the decoupling between spin and charge degrees of freedom can support the appearance a partly gapped DQCP which in the specific BOW-AF case is characterized by $\Delta_{c}\neq 0$ while having long-range order of $C_{P}^{c}(r)$.
@@ III.2.3 AF-CDW phase transition
Finally, we study the AF-CDW phase transition. As already pointed out at the level of field theory, in order to find such transition to be continuous, both gaps have to vanish at the transition point. This aspect clearly rules out the possible presence of a partly gapped DQCPs while, in principle, it still allows for standard fully gapless DQCPs where all the correlators decay algebraically. As already discussed, the presence of such a DQCPs would require a fine tuning of the microscopic parameters in order to have a simultaneous closing of both the spin and the charge gap. In this regard, our numerical analysis finds that the AF-CDW transition is always first order for all values of $U/t$. In Fig. 7 we prove this for the case of $U/t=8$. As can be seen, we find indeed that both the order parameters ${\cal O}_{CDW}$ ${\cal O}_{AF}$ and the Luttinger parameters display a discontinuity at the transition point, see Fig. 7(a) and (b) respectively. Moreover, while the former quantities are not affected by different $\chi$ values, the latter always show size dependence associated to almost vanishing values.
@@ IV Conclusions
We unveiled a new type of deconfined quantum critical points. Contrary to previously known examples where they are associated to fully gapless excitations, we have shown that the phenomenon of spin and charge separation can generate partly gapped deconfined quantum critical points. Specifically, by means of analytical and numerical treatments, we have been able to prove that specific phase transitions connecting locally ordered fermionic phases are not captured by the Landau-Ginzburg-Wilson symmetry-breaking paradigm. Notably, at these phase transitions we find that only one gap closes while the other remains finite. Therefore, long-range order of specific nonlocal correlation functions occurs. In addition, we calculated the critical exponents governing how the local order parameters vanish. In such a way, we demonstrated that these transitions are one-dimensional realizations of deconfined quantum critical points with the intriguing feature of being only partly gapped. In conclusion, we also point out that at least one of these phase transitions, the one between the bond-order-wave and charge-density-wave, has been predicted to characterize the phase diagram of ultracold dipolar atoms in optical lattice [93, 78]. Here, the presence of partially gapped deconfined quantum critical points can be probed by measuring the defect production [94, 95, 96] predicted by the Kibble-Zurek mechanism [97, 98, 99, 100, 101, 102] in continuous transitions. In addition, quantum gas microscopy [103] and noise correlators measurements [104] allow for an accurate detection of local density [105] and spin [106] orderings and therefore for an accurate probing of nonlocal order parameters and locally ordered phases. As consequence, we believe that our results not only unveil a novel interesting type of deconfined quantum critical points but also pave the way towards their experimental detection, which represents a fundamental challenge solid state platforms.  [107, 108, 109, 110].